# AAPO: Enhance the Reasoning Capabilities of LLMs with Advantage Momentum

### v1

@@ 5.2 Understanding the effectiveness of AAPO
$\mathcal{J}_{\mathrm{AAPO}}(\theta)$ $={\frac{1}{\sum_{i=1}^{G}\left|o_{i}\right|}\sum_{i=1}^{G}\sum_{t=1}^{\left|o_{i}\right|}\left(-\log\pi_{\theta}\left(o_{i,t}\mid q,o_{i,<t}\right)\hat{A}_{i,t}^{*}\right)}$
$\approx{\frac{1}{\sum_{i=1}^{G}\left|o_{i}\right|}\sum_{i=1}^{G}\sum_{t=1}^{\left|o_{i}\right|}\left(-\log\pi_{\theta}\left(o_{i,t}\mid q,o_{i,<t}\right)\cdot\operatorname{clip}(r_{\theta_{i}}-r_{ref_{i}},\delta_{\mathrm{low}},\delta_{\mathrm{high}})\right)}$
The optimization exhibits slight oscillations near the global optimum, avoiding large gradient updates.
6 Experiments
6.1 Experimental setup
Most recent RL algorithms [36, 48, 18, 25] struggle to make LLMs perform better at solving mathematical problems, which requires the model to think in the CoT [43] format before deciding the final answer. We choose open-rs [10] as our training dataset for DeepSeek-R1-Distill-Qwen-1.5B base model, because the data in this dataset cover various types and difficulty levels of mathematical problems that are highly representative. To further provide a fair and rigorous evaluation of the effectiveness of our proposed AAPO, Qwen2.5-Math-7B model is chosen as the base model and subsequently trained on more challenging simplelr_qwen_level3to5 dataset [44].
In our experiment setting, we set the clip parameters $\delta_{\mathrm{low}}$ and $\delta_{\mathrm{high}}$ to be -0.2 and 0.28, respectively. We train all base models under the AAPO following the training process depicted in Algorithm 1. All rule-based reward functions adopted in our experiments are simple and straightforward. More training details about rule-based reward functions and the system prompt are provided in Appendix A. To evaluate the extent to which our proposed AAPO algorithm can enhance the reasoning capabilities of the model, we select AIME24, MATH-500 [17, 23], AMC23, Minerva [22], and OlympiadBench [20] as evaluation benchmarks. Our evaluation framework utilizes a well-established and community-vetted codebase, maintaining consistency with widely adopted implementations.
Figure 2: A comparison of the evaluation results at each epoch during the training of the Qwen2.5-Math-7B model on simplelr_qwen_level3to5, w/o and w/ the clip operation.
6.2 Results
As shown in Table 1, taking model DeepSeek-R1-Distill-Qwen-1.5B as our base model, the application of our proposed AAPO enables the base model to achieve the SOTA performance on Minerva, the second-best performance on MATH-500, AMC23 and OlympiadBench. When averaging scores across all benchmarks, the resulting AAPO-1.5B model achieves an overall SOTA performance. By direct comparison under the same training data, AAPO-1.5B achieves improvements of 2.7% and 2.0% over GRPO-1.5B and GPG-1.5B, respectively. It is worth noting that our AAPO-1.5B achieves performance superior to Still-3-1.5B-Preview [39], despite the fact that Still-3-1.5B-Preview benefits from a larger training dataset that consists of long CoT reasoning data distilled from the DeepSeek-R1 [14] model and performs RL training with specified reward strategies. After employing our proposed AAPO on the Qwen2.5-Math-7B model, AAPO-7B achieves the overall SOTA performance compared to other methods. AAPO-7B achieves improvements of 12.2%, 10.3%, 4.2% over SimpleRL-Zero-7B [44], GPG-7B [8], and Oat-Zero-7B [25], respectively, under identical training data. Furthermore, AAPO-7B also outperforms other models such as Eurus-2-7B-PRIME [9] and OpenReasoner-Zero-7B [19], despite the fact that these baselines are trained with more high-quality data or data distilled from the DeepSeek-R1 model. This suggests that the effectiveness can be mainly attributed to the design of our proposed AAPO rather than to the scale or quality of the training data.
Model AIME24 MATH-500 AMC23 Minerva OlympiadBench Avg.
7B Models
Qwen2.5-Math-7B [46] 13.3 57.6 45.0 14.7 23.7 30.9
SimpleRL-Zero-7B [44] 30.0 77.4 57.5 30.5 38.1 46.7
GPG-7B [8] 23.3 80.2 55.0 36.0 42.8 47.5
OpenReasoner-Zero-7B [19] 20.0 80.8 65.0 29.4 46.2 48.3
Eurus-2-7B-PRIME [9] 16.7 81.8 65.0 37.5 44.6 49.1
Oat-Zero-7B [25] 30.0 81.2 65.0 34.9 43.4 50.3
rStar-Math-7B [12] 26.7 78.4 47.5 47.1
AAPO-7B (Ours) 30.0 82.4 70.0 35.3 44.3 52.4
1.5B Models
DeepSeek-R1-Distill-Qwen-1.5B 28.8 82.8 62.9 26.5 43.3 48.9
GRPO-1.5B [10] 26.7 86.2 82.5 27.6 52.6 55.2
GPG-1.5B [8] 36.7 83.4 75.0 29.8 53.2 55.6
Still-3-1.5B-Preview [4] 40.0 85.5 72.5 30.5 53.9 56.5
AAPO-1.5B (Ours) 33.3 86.0 80.0 30.9 53.3 56.7
Table 1: Zero-shot pass@1 performance on mathematical reasoning benchmarks. represents reproduced results with our best effort. Dashes (–) denote unavailable official scores. Bold and Underline indicate the best and the second-best performance in the corresponding category, respectively.
6.3 Ablation study
Clip operation To investigate the contribution of the clip operation to AAPO, we conduct additional ablation studies by removing the clip operation on both the 1.5B and 7B models. The results presented in Table 2 indicate that the performance of the optimized model without clip is inferior to that with the clip, suggesting that incorporating the clip operation effectively contributes to further improvements in optimization results. As illustrated in Figure 2, the optimization process becomes more stable with the incorporation of the clip operation compared to the optimization process without it. The resulting AAPO-7B exhibits better performance on the benchmarks when the clip operation is adopted.
Model AIME24 MATH-500 AMC23 Minerva OlympiadBench Avg.
AAPO-1.5B 33.3 86.0 80.0 30.9 53.3 56.7
AAPO-1.5B w/o clip 33.3 85.0 82.5 29.0 53.3 56.6
AAPO-7B 30.0 82.4 70.0 35.3 44.3 52.4
AAPO-7B w/o clip 30.0 79.6 70.0 34.9 42.5 51.2
Training process analysis  As illustrated by the training curves in Figure 3, AAPO achieves optimization performance comparable to that of GPG, while exhibiting further improvements during the later steps of RL training. In the Format Reward and Reward figures, AAPO consistently attains higher reward than GPG in the later steps of the training. In addition, three Reward figures demonstrate significantly reduced fluctuations during the training process. AAPO also outperforms GRPO. Moreover, in the Response Length figure, our proposed AAPO demonstrates superior optimization, indicating higher training efficiency. The final results presented in Table 1 demonstrate that our proposed AAPO achieves superior performance across five mathematical reasoning benchmarks.
Figure 3: Training process of DeepSeek-R1-Distill-Qwen-1.5B on open-rs [10] utilizing our proposed AAPO algorithm. Compared to GPG [8] and GRPO [36], AAPO demonstrates better stability during training and achieves superior performance in the final results as shown in Table 1.
7 Conclusion
In this paper, we conduct an in-depth analysis of the limitations inherent in the group relative advantage estimation method used by mainstream RL algorithms, such as GRPO, which would lead to optimization issues such as zero gradient and gradient ascent. To address these issues, we propose a novel RL algorithm Advantage-Augmented Policy Optimization (AAPO). By augmenting the group relative advantage estimation method with advantage momentum, our method effectively improves policy optimization performance in experimental benchmarks. Experimental results demonstrate that AAPO achieves the overall superior performance across several mathematical reasoning benchmarks.
Limitations  Although AAPO demonstrates superior performance, the main limitation of our work lies in the substantial computational resources and the long training time required.
Appendix
Appendix A Experiment details
A.1 Reward rules
Format Reward  To encourage adherence to structured reasoning, we adopt a binary format reward $R_{format}(o)\in\{0,1\}$, which assigns a reward of 1 if the model response $o$ conforms to the expected template by containing the delimiter sequence "\$\backslash\mathrm{n}</\text{think}>\backslash\mathrm{n}\text{"}$, and 0 otherwise.
@@ A.1 Reward rules
Cosine Scaled Reward  We adopt the $R_{cosine\_scaled\_accuracy}\in\{0,1\}$ as expressed in Equation (10), which encourages correct outputs with shorter lengths and penalizes incorrect outputs with reduced severity as their length increases, following a cosine annealing schedule in Equation (11).
@@
$R_{\text{correct}}(l)&=\alpha_{\min}^{c}+\frac{1}{2}(\alpha_{\max}^{c}-\alpha_{\min}^{c})\left[1+\cos\left(\pi\frac{l}{L}\right)\right],\\ R_{\text{wrong}}(l)&=\alpha_{\max}^{w}+\frac{1}{2}(\alpha_{\min}^{w}-\alpha_{\max}^{w})\left[1+\cos\left(\pi\frac{l}{L}\right)\right],\end{aligned}.$ (11)
@@ A.2 Training setup
Training DeepSeek-R1-Distill-Qwen-1.5B  For training DeepSeek-R1-Distill-Qwen-1.5B on open-rs [10] experiment, we adopt the $R_{format}$ and $R_{cosine\_scaled\_accuracy}$ reward functions, with respective weights of 1 and 2. We directly perform our AAPO training from the base model without any SFT with the group size of 6 and the batch size of 6 on 2 Nvidia A800 GPUs with 80G VRAM. The system prompt adopted in the RL training is provided in Figure 4.
Figure 4: System prompt for training DeepSeek-R1-Distill-Qwen-1.5B base model.
@@
Training Qwen2.5-Math-7B  For training Qwen2.5-Math-7B [46] on simplelr_qwen_level3to5 [44], we adopt the $R_{accuracy}$ reward function with a weight of 1. We directly perform our AAPO training without any SFT with the group size of 8 and the batch size of 16 on 8 Nvidia A800 GPUs with 80G VRAM. The system prompt adopted in the RL training is provided in Figure 5.
Figure 5: System prompt for training Qwen2.5-Math-7B base model.
A.3 Evaluation setup
@@
All the reproduced results in Table 1 are evaluated using the lighteval [16] framework with vllm backend proposed by Hugging Face. Our evaluation experiments on all benchmarks are conducted on a single Nvidia A800 GPU with 80G VRAM, with system prompt provided in Figure 6.
Figure 6: System prompt for evaluation on all benchmarks.
Appendix B More analysis on experiment results
@@
We plot the loss curves of AAPO and GPG [8] during training DeepSeek-R1-Distill-Qwen-1.5B on open-rs [10] in Figure 7. As observed, the loss values for AAPO remain predominantly positive throughout the training process. This indicates that AAPO primarily optimizes the policy by encouraging diverse responses, assigning different gradient magnitudes according to the advantage estimated by $\hat{A}_{i,t}^{*}$, thus leading to performance improvements. In contrast, GPG exhibits mostly negative loss values, suggesting that it focuses on suppressing suboptimal responses as its main optimization strategy. These results of the analysis imply that models trained with our proposed AAPO demonstrate stronger generalization, resulting in superior performance as shown in Table 1.
@@
Figure 7: A comparative analysis of training loss between AAPO and GPG.

### v2

@@ Appendix B Proof
$\mathbb{E}\!\left[\nabla\mathcal{L}_{\mathcal{G}}(\theta_{k})\right]=\nabla\mathcal{L}(\theta_{k}),\quad\mathbb{E}\!\left[\left\|\nabla\mathcal{L}_{\mathcal{G}}(\theta_{k})-\nabla\mathcal{L}(\theta_{k})\right\|^{2}\right]\leq\frac{\sigma^{2}}{N_{\mathcal{G}}}.$
Consequently (by variance decomposition),
$\mathbb{E}\!\left[\left\|\nabla\mathcal{L}_{\mathcal{G}}(\theta_{k})\right\|^{2}\right]\leq\left\|\nabla\mathcal{L}(\theta_{k})\right\|^{2}+\frac{\sigma^{2}}{N_{\mathcal{G}}}.$
Taking total expectation and noting $\mathbb{E}[\mathcal{L}_{\mathcal{G}}(\theta)]=\mathcal{L}(\theta)$, we obtain
$\mathbb{E}\!\left[\mathcal{L}(\theta_{k+1})\right]\leq\mathbb{E}\!\left[\mathcal{L}(\theta_{k})\right]-\tfrac{\eta_{k}}{2}\,\mathbb{E}\!\left[\left\|\nabla\mathcal{L}(\theta_{k})\right\|^{2}\right]+\tfrac{L}{2}\,\tfrac{\eta_{k}^{2}\sigma^{2}}{N_{\mathcal{G}}}.$
Since $\mathcal{L}(\theta)$ is lower bounded (Theorem 1), applying the Robbins–Siegmund theorem (Robbins & Siegmund, 1971) gives
$\sum_{k=0}^{\infty}\eta_{k}\,\mathbb{E}\!\left[\left\|\nabla\mathcal{L}(\theta_{k})\right\|^{2}\right]<\infty.$
Given $\sum_{k}\eta_{k}=\infty$, it follows that
$\liminf_{k\to\infty}\mathbb{E}\!\left[\left\|\nabla\mathcal{L}(\theta_{k})\right\|^{2}\right]=0.$
If $\eta_{k}\equiv\eta\leq 1/L$ is fixed, the residual is of order $\mathcal{O}(\eta)$ (from smoothness) and $\mathcal{O}(1/N_{\mathcal{G}})$ (from gradient variance). Hence
$\limsup_{K\to\infty}\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}\!\left[\left\|\nabla\mathcal{L}(\theta_{k})\right\|^{2}\right]\;\lesssim\;\mathcal{O}(\eta)+\mathcal{O}\!\left(\tfrac{1}{N_{\mathcal{G}}}\right).$
With diminishing step sizes, the algorithm converges to a stationary point of the expected objective; with a small constant step size, it converges to a neighborhood of stationarity whose size depends on $\eta$ and the group size. The advantage momentum does not affect the asymptotic rates, as it only changes the constant $B$ (and hence $L=BL_{0}$). ∎
Appendix C Experiment details
C.1 Reward rules
[…]
@@ C.1 Reward rules
Cosine Scaled Reward  We adopt the $R_{cosine\_scaled\_accuracy}\in\{0,1\}$ as expressed in equation 9, which encourages correct outputs with shorter lengths and penalizes incorrect outputs with reduced severity as their length increases, following a cosine annealing schedule in equation 10.
@@
$\begin{aligned} R_{\text{correct}}(l)&=\alpha_{\min}^{c}+\frac{1}{2}(\alpha_{\max}^{c}-\alpha_{\min}^{c})\left[1+\cos\left(\pi\frac{l}{L}\right)\right],\\ R_{\text{wrong}}(l)&=\alpha_{\max}^{w}+\frac{1}{2}(\alpha_{\min}^{w}-\alpha_{\max}^{w})\left[1+\cos\left(\pi\frac{l}{L}\right)\right],\end{aligned}.$ (10)
@@ C.2 Training setup
Training DeepSeek-R1-Distill-Qwen-1.5B  For training DeepSeek-R1-Distill-Qwen-1.5B on open-rs experiment (Dang & Ngo, 2025), we adopt the $R_{format}$ and $R_{cosine\_scaled\_accuracy}$ reward functions, with respective weights of 1 and 2. We directly perform our AAPO training from the base model without any SFT with the group size of 6 and the per_device_batch_size of 6 with gradient_accumulation_steps of 4 on 2 Nvidia A800 GPUs with 80G VRAM. The system prompt adopted in the RL training is provided below.
System prompt for training DeepSeek-R1-Distill-Qwen-1.5B base model You are a helpful AI Assistant, designed to provided well-reasoned anddetailed responses. You FIRST think about the reasoning process as an internal monologue and then provide the user with the answer. The reasoning process MUST BE enclosed within <think>and </think>tags.
@@
Training Qwen2.5-Math-7B  For training Qwen2.5-Math-7B (Yang et al., 2024) on simplelr_qwen_level3to5 (Zeng et al., 2025), we adopt the $R_{accuracy}$ reward function with a weight of 1. We directly perform our AAPO training without any SFT with the group size of 8 and the per_device_batch_size 8 with gradient_accumulation_steps of 4 on 8 Nvidia A800 GPUs with 80G VRAM. The system prompt adopted in the RL training is provided below.
[…]
Training Llama series models  For training Llama-3.2-1B-Insrtuct and Llama-3.2-3B-Instruct models, we adopt the simplelr_able_level3to5 dataset (Zeng et al., 2025) and the $R_{accuracy}$ reward function with a weight of 1. We directly perform our AAPO training without any SFT with the group size of 8 and the per_device_batch_size of 16 on 4 Nvidia A800 GPUs with 80G VRAM. There is no extra system prompt added when training Llama series models.
C.3 Evaluation setup
@@
Evaluation of Qwen series models  All the reproduced results in Table 1 are evaluated using the lighteval framework (Habib et al., 2023) with vllm backend proposed by Hugging Face. Our evaluation experiments on all benchmarks are conducted on a single Nvidia A800 GPU with 80G VRAM, with system prompt provided in below.
[…]
Evaluation of Llama series models  All the reproduced results in Table 2 are evaluated using vllm backend for generation and the evaluation script provided by Zeng et al. (2025). Our evaluation experiments on all benchmarks are conducted on 4 Nvidia A800 GPUs with 80G VRAM.
Appendix D More analysis on experiment results
@@
We plot the loss curves of AAPO and GPG (Chu et al., 2025) during training DeepSeek-R1-Distill-Qwen-1.5B on open-rs (Dang & Ngo, 2025) in Figure 3. As observed, the loss values for AAPO remain predominantly positive throughout the training process. This indicates that AAPO primarily optimizes the policy by encouraging diverse responses better than the reference, assigning different gradient magnitudes according to the advantage estimated by $\hat{A}_{i,t}^{*}$, thus leading to performance improvements. In contrast, GPG exhibits mostly negative loss values, suggesting that it focuses on suppressing suboptimal responses as its main optimization strategy. It can be clearly seen from the Figure 3 that the training process of AAPO is more stable than GRPO (Shao et al., 2024). These results of the analysis imply that models trained with our proposed AAPO demonstrate stronger generalization, resulting in superior performance as shown in Table 1.
@@
Figure 3: A comparative analysis of training loss between AAPO and GPG, AAPO and GRPO.
Appendix E Extra comparison with original results
Here we present the original results reported in the corresponding papers for Qwen series models.
Model AIME24 MATH-500 AMC23 Minerva OlympiadBench Avg.
Qwen 1.5B Models
DeepSeek-R1-Distill-Qwen-1.5B 28.9 83.9
GRPO-1.5B (Dang & Ngo, 2025) 46.7 84.4 72.5 26.8 51.3 56.3
GPG-1.5B (Chu et al., 2025) 33.3 85.0 80.0 26.8 52.4 55.5
Still-3-1.5B-Preview (Chen et al., 2025) 39.3 85.5
[…]
Qwen 7B Models
Qwen2.5-Math-7B (Yang et al., 2024) 55.4
SimpleRL-Zero-7B (Zeng et al., 2025) 20.0 78.2 62.5 38.6 40.4 47.9
GPG-7B(Chu et al., 2025) 33.3 80.0 65.0 34.2 42.4 51.0
OpenReasoner-Zero-7B (Hu et al., 2025)
Eurus-2-7B-PRIME (Cui et al., 2025) 20.0 78.2 50.6 39.3 40.3 45.7
Oat-Zero-7B(Liu et al., 2025) 43.3 80.0 62.7 30.1 41.0 51.4
[…]
Table 3: Zero-shot pass@1 performance on mathematical reasoning benchmarks. All reported results in this table are directly adopted from the corresponding papers. Dashes (–) denote unavailable official score.